What about in the opposite scenario? If the central bank withdraws $5000 from circulation, in a population with a currency ratio of 0.05 and a reserve ratio of 0.1, what is the final effect on the money supply?
Answer #1

Money multiplier = (1+c)/(c+r) = (1+0.05)/(0.05+0.1)= 7

Money supply decreases by 7*5000= 35000